# Null projections and noncommutative function theory in operator algebras

### v1

@@ 7.1. Von Neumann algebraic preliminaries
We being with some general facts about von Neumann algebras that will be needed throughout the remainder of the paper.
Given a von Neumann algebra $M$, we let $M_{*}\subset M^{*}$ denote the norm-closed subspace of weak-*** continuous linear functionals on $M$. The following is actually a general principle for dual Banach spaces that is well known in certain quarters. It is a necessary and sufficient condition for a kind of ‘Kaplansky density’.
@@
Let $M$ be a von Neumann algebra. Let $A\subset M$ be a subspace and let $W\subset M$ denote its weak-*** closure. Let $\rho:M_{*}\to A^{*}$ denote the restriction map. Then, $\rho$ has closed range if and only if there is a constant $r>0$ such that the closed unit ball of $W$ is contained in the weak-*** closure of $\{a\in A:\|a\|\leq r\}$.
@@
By the convex separation theorem, we conclude that the weak-*** closure of $\{a\in A:\|a\|\leq r\}$ contains the closed unit ball of $W$.
Conversely, assume that there is $r>0$ such that the closed unit ball of $W$ is contained in the weak-*** closure of $\{a\in A:\|a\|\leq r\}$. Let $\omega\in M_{*}$. It follows that
@@
There is a central projection $z\in M^{**}$ such that $M_{*}=M^{*}z$ [takesaki2002, Theorem III.2.14]. An important fact is that there is a normal ***-isomorphism $\Omega:M\to M^{**}z$ such that $\Omega(a)=az$ for every $a\in M$. Next, let $\varphi$ be a normal state on $M$, and let $\widehat{\varphi}:M^{**}\to{\mathbb{C}}$ denote its unique normal extension. Consider the weak-*** closed left ideals
@@
More generally, we can describe $M_{*}$ in terms of absolute continuity, as follows. We denote by $S_{n}(M)$ the set of normal states on $M$. Assume now that $M$ contains a weak-*** dense unital $\mathrm{C}^{*}$-subalgebra $B$. Let $\rho:M^{*}\to B^{*}$ denote the restriction map. By the Kaplansky density theorem we see that $\rho$ is isometric on $M_{*}$. In particular, this means that the sets $(M_{*})|_{B}=\rho(M_{*})$ and $\rho(S_{n}(M))$ are norm-closed. Put $\Delta=\rho(S_{n}(M))$; this is the norm-closed convex set in $B^{*}$ consisting of those elements of the form $\omega|_{B}$, where $\omega$ is a normal state on $M$. We now record a simple fact.
@@
We have that $*\operatorname{AC}(S_{n}(M)))=M_{*}$ and $\operatorname{AC}(\Delta)=(M_{*})|_{B}$.
@@
Given $\omega\in M_{*}$ and $a\in M$, we have that $*\omega(\cdot a)\in M_{*}$. Because both $M_{*}$ and $(M_{*})|_{B}$ are norm-closed, we then find $\operatorname{AC}(S_{n}(M))\subset M_{*}$ and $\operatorname{AC}(\Delta)\subset(M_{*})|_{B}$ by virtue of [CT2023Henkin, Theorem 2.7]. Conversely, let $\varphi\in M_{*}$ with $\|\varphi\|=1$. We have $\varphi(z\cdot)=\varphi$. Now, there is a partial isometry $v\in M$ such that $|\varphi|=\varphi(\cdot v)$. It follows that $|\varphi|(z\cdot)=|\varphi|$, whence $|\varphi|\in S_{n}(M)$ and $\varphi\in\operatorname{AC}(S_{n}(M))$. We conclude that $M_{*}\subset\operatorname{AC}(S_{n}(M))$, which in turn plainly implies $(M_{*})|_{B}\subset\operatorname{AC}(\Delta)$. ∎
@@ 7.2. Lebesgue decompositions
Next, let $M$ be a von Neumann algebra containing $B$ as a weak-*** dense unital $\mathrm{C}^{*}$-subalgebra. We say that $A$ admits a Lebesgue decomposition relative to $M$ if $(M_{*})|_{A}$ is norm-closed and there is a (necessarily unique) subspace $\Sigma\subset A^{*}$ such that $A^{*}=(M_{*})|_{A}\oplus_{1}\Sigma$.
It follows from [CH2022, Lemma 3.4] that $B$ itself always admits a Lebesgue decomposition relative to $M$. By contrast, the existence of such a decomposition for $A$ is somewhat rare (see [CH2022] for more detail). Combining Lemma 7.3 with [CH2022, Theorem 2.4], we see that if $A$ has the F.& M. Riesz property in $B$ with respect to the convex set of restrictions to $B$ of the normal states on $M$, then $A$ admits a Lebesgue decomposition relative to $M$ that is ‘compatible’ (in a precise sense) with that of $B$. We can now give a different sufficient condition, where we don’t insist on this compatibility condition.
@@
Let $M$ be a von Neumann algebra. Let $B\subset M$ be a weak-*** dense unital $\mathrm{C}^{*}$-subalgebra, and let $A\subset B$ be a unital commutative subalgebra. Let
@@
so that $\varphi\in(A^{**}(1-r_{\Gamma}))_{\perp}$. Conversely let $\varphi\in A^{*}$ such that $\varphi((1-r_{\Gamma})\cdot)=0$. By the Hahn–Banach theorem, there is $\psi\in B^{*}$ such that $\psi|_{A}=\varphi$. Since $r_{\Gamma}\in A^{\perp\perp}$, this implies that $\psi(r_{\Gamma}\cdot)|_{A}=\varphi(r_{\Gamma}\cdot)=\varphi$. Put $*\psi^{\prime}=\psi(r_{\Gamma}\cdot)\in B^{*}$, which then lies in $\operatorname{AC}(\Gamma)=X$. Hence, there is $\omega\in M_{*}$ such that $\omega|_{A}=\psi^{\prime}|_{A}=\varphi$, so that $\varphi\in(M_{*})|_{A}$ as desired. ∎
@@
Let $d\geq 1$ be an integer, let ${\mathbb{B}}_{d}\subset{\mathbb{C}}_{d}$ be the open unit ball, and let ${\mathbb{S}}_{d}$ denote its topological boundary, the unit sphere. Let $\sigma$ denote surface measure on ${\mathbb{S}}_{d}$. Let $M=L^{\infty}({\mathbb{S}}_{d},\sigma),B=\mathrm{C}({\mathbb{S}}_{d})$ and let $A\subset B$ denote the ball algebra, i.e. the norm-closure of the polynomials. It is well known that the weak-*** closure of $A$ in $M$ is simply $H^{\infty}({\mathbb{B}}_{d})$, by making the usual identification via radial boundary values. Further, given $f\in H^{\infty}({\mathbb{B}}_{d})$ and $10<r<$ 1, we define $f_{r}:\overline{{\mathbb{B}}_{d}}\to{\mathbb{C}}$ as $f_{r}(z)=f(rz)$. Then, $f_{r}\in A$, $\|f_{r}\|\leq\|f\|$ and the net $(f_{r})$ converges to $f$ in the weak-*** topology of $M$. We infer that the set
@@
Next, we note that $B$ admits a Lebesgue decomposition relative to $M$ by [CH2022, Lemma 3.4]. Nevertheless, the Lebesgue decompositions of $A$ and $B$ relative to $M$ are not compatible, by [CH2022, Theorem 4.8 and Proposition 6.3]. ∎
@@
Next, recall that ${\mathcal{T}}({\mathcal{H}})$ contains the ideal of compact operators on ${\mathcal{H}}$, whence it admits a Lebesgue decomposition relative to $B({\mathcal{H}})$ by [CH2022, Lemma 3.4]. Nevertheless, the Lebesgue decompositions of ${\mathcal{A}}({\mathcal{H}})$ and ${\mathcal{T}}({\mathcal{H}})$ relative to $B({\mathcal{H}})$ are not compatible, by [CH2022, Theorem 4.8 and Proposition 6.1]. ∎
We emphasize again that the incompatibility of the Lebesgue decompositions in the previous two examples shows in particular that Corollary 7.4 is applicable in some cases where [CH2022, Theorem 2.4] is not.
@@
This set will play an analogous role to the set of functionals whose restrictions to $A$ are “Henkin” (see e.g. Section 8.1 below). We now show that the F.& M. Riesz property implies a certain rigidity property for left bands containing $\Delta$. Its relation to the the classical Cole-Range theorem [rudin2008, Theorem 9.6.1] will be thoroughly explained at the end of this section.
@@
    $A$ has the strong F.& M. Riesz property in $B$ with respect to $\Delta$.
@@
Let $\lambda$ denote arclength measure on the unit circle ${\mathbb{T}}\subset{\mathbb{C}}$. By means of radial boundary values, we may embed the algebra $H^{\infty}({\mathbb{D}})$ of bounded holomorphic functions on the open unit disc as a unital norm-closed subalgebra of $L^{\infty}({\mathbb{T}},\lambda)$. Choosing $\Delta=\{\lambda\}$, Lemma 7.2 implies that $\operatorname{AC}(\Delta)$ coincides with the weak-*** continuous linear functionals on $L^{\infty}({\mathbb{T}},\lambda)$. Hence, $E(\Delta)$ coincides with the set $X$ appearing in Corollary 7.4.
Consider now the Douglas algebra $H^{\infty}({\mathbb{D}})+\mathrm{C}({\mathbb{T}})\subset L^{\infty}({\mathbb{T}},\lambda)$, which is known to be proper and norm-closed (see for instance [douglas1998, Chapter 6]). Hence, there is a non-zero bounded linear functional $\varphi$ on $L^{\infty}({\mathbb{T}},\lambda)$ that annihilates $H^{\infty}({\mathbb{D}})+\mathrm{C}({\mathbb{T}})$. Since $\mathrm{C}({\mathbb{T}})$ is weak-*** dense in $L^{\infty}({\mathbb{T}},\lambda)$, we see that $\varphi$ is not weak-*** continuous, so that $H^{\infty}({\mathbb{D}})$ does not have the strong F.& M. Riesz property in $L^{\infty}({\mathbb{T}},\lambda)$ with respect to $\Delta$. On the other hand, it is known $H^{\infty}({\mathbb{D}})$ that has the F.& M. Riesz property in $L^{\infty}({\mathbb{T}},\lambda)$ with respect to $\Delta$ (see [ando1978]). We may thus use Corollary 8.2 to see that the set $E(\Delta)$ is not a left band. ∎
@@ 8.1. Henkin functionals and the Cole-Range phenomenon
Let $M$ be a von Neumann algebra. Let $B\subset M$ be a unital $\mathrm{C}^{*}$-subalgebra that is weak-*** dense in $M$. Let $A\subset B$ be a unital subalgebra. A bounded linear functional $\varphi:A\to{\mathbb{C}}$ is said to be Henkin relative to $M$ if it extends weak-*** continuously to the weak-*** closure of $A$ in $M$. This notion is in agreement with our terminology used in Section 5. Equivalently, $\varphi$ is Henkin if it extends weak-*** continuously to $M$ [rudin1991FA, Theorem 3.6], so the space of Henkin functionals is simply $(M_{*})|_{A}$.
@@
Clearly, integration against $\sigma$ is in $\Omega$ by Cauchy’s formula. Furthermore, $\Omega\subset X$. Since Henkin functionals on $A$ admit a weak-*** continuous extension to $M$, it follows that $X\subset A^{\perp}+\operatorname{AC}(\sigma)\subset A^{\perp}+\operatorname{AC}(\Omega)=E(\Omega)$.
@@ 9. The Amar–Lederer property
Consider $H^{\infty}({\mathbb{D}})\subset L^{\infty}({\mathbb{T}},\lambda)$ as in Example 5. Let $X$ denote the character space of $L^{\infty}({\mathbb{T}},\lambda)$. Then, via the Gelfand transform, we know that $L^{\infty}({\mathbb{T}},\lambda)\cong\mathrm{C}(X)$. Let $A\subset\mathrm{C}(X)$ denote the unital norm-closed subalgebra corresponding to $H^{\infty}({\mathbb{D}})$ under this identification. Let $\tau$ denote the state on $\mathrm{C}(X)$ corresponding to integration against $\lambda$ on $L^{\infty}({\mathbb{T}},\lambda)$. Let $\Delta=\{\tau\}$. In this case, $A$ has the Amar–Lederer property in $B$ with respect to $\Delta$ by a classical result of Amar and Lederer [AL1971]. Of course understanding this requires also the dictionary between closed sets and closed projections, and peak sets and closed projections in $A^{\perp\perp}$, that we have discussed earlier. On the other hand, properties (iv) and (v) in Theorem 6.5 fail. To see this note that the latter properties, in the specific case of this example, imply that closed ‘Lebesgue-null’ sets are generalized peak sets. Yet [pelczynski1977, Proposition 6.3] gives a closed ‘Lebesgue-null’ $G_{\delta}$ set in this context which is not a peak set. Hence, it is not a generalized peak set, since $G_{\delta}$ generalized peak sets are peak sets [gamelin1969, Lemma II.12.1].

### v2

@@ 7.1. Von Neumann algebraic preliminaries
We begin with some general facts about von Neumann algebras that will be needed throughout the remainder of the paper.
Given a von Neumann algebra $M$, we let $M_{*}\subset M^{*}$ denote the norm-closed subspace of weak-* continuous linear functionals on $M$. The following is actually a general principle for dual Banach spaces that is well known in certain quarters. It is a necessary and sufficient condition for a kind of ‘Kaplansky density’.
@@
Let $M$ be a von Neumann algebra. Let $A\subset M$ be a subspace and let $W\subset M$ denote its weak-* closure. Let $\rho:M_{*}\to A^{*}$ denote the restriction map. Then, $\rho$ has closed range if and only if there is a constant $r>0$ such that the closed unit ball of $W$ is contained in the weak-* closure of $\{a\in A:\|a\|\leq r\}$.
@@
By the convex separation theorem, we conclude that the weak-* closure of $\{a\in A:\|a\|\leq r\}$ contains the closed unit ball of $W$.
Conversely, assume that there is $r>0$ such that the closed unit ball of $W$ is contained in the weak-* closure of $\{a\in A:\|a\|\leq r\}$. Let $\omega\in M_{*}$. It follows that
@@
There is a central projection $z\in M^{**}$ such that $M_{*}=M^{*}z$ [takesaki2002, Theorem III.2.14]. An important fact is that there is a normal *-isomorphism $\Omega:M\to M^{**}z$ such that $\Omega(a)=az$ for every $a\in M$. Next, let $\varphi$ be a normal state on $M$, and let $\widehat{\varphi}:M^{**}\to{\mathbb{C}}$ denote its unique normal extension. Consider the weak-* closed left ideals
@@
More generally, we can describe $M_{*}$ in terms of absolute continuity, as follows. We denote by $S_{n}(M)$ the set of normal states on $M$. Assume now that $M$ contains a weak-* dense unital $\mathrm{C}^{*}$-subalgebra $B$. Let $\rho:M^{*}\to B^{*}$ denote the restriction map. By the Kaplansky density theorem we see that $\rho$ is isometric on $M_{*}$. In particular, this means that the sets $(M_{*})|_{B}=\rho(M_{*})$ and $\rho(S_{n}(M))$ are norm-closed. Put $\Delta=\rho(S_{n}(M))$; this is the norm-closed convex set in $B^{*}$ consisting of those elements of the form $\omega|_{B}$, where $\omega$ is a normal state on $M$. We now record a simple fact.
@@
We have that $\operatorname{AC}(S_{n}(M)))=M_{*}$ and $\operatorname{AC}(\Delta)=(M_{*})|_{B}$.
@@
Given $\omega\in M_{*}$ and $a\in M$, we have that $\omega(\cdot a)\in M_{*}$. Because both $M_{*}$ and $(M_{*})|_{B}$ are norm-closed, we then find $\operatorname{AC}(S_{n}(M))\subset M_{*}$ and $\operatorname{AC}(\Delta)\subset(M_{*})|_{B}$ by virtue of [CT2023Henkin, Theorem 2.7]. Conversely, let $\varphi\in M_{*}$ with $\|\varphi\|=1$. We have $\varphi(z\cdot)=\varphi$. Now, there is a partial isometry $v\in M$ such that $|\varphi|=\varphi(\cdot v)$. It follows that $|\varphi|(z\cdot)=|\varphi|$, whence $|\varphi|\in S_{n}(M)$ and $\varphi\in\operatorname{AC}(S_{n}(M))$. We conclude that $M_{*}\subset\operatorname{AC}(S_{n}(M))$, which in turn plainly implies $(M_{*})|_{B}\subset\operatorname{AC}(\Delta)$. ∎
@@ 7.2. Lebesgue decompositions
Next, let $M$ be a von Neumann algebra containing $B$ as a weak-* dense unital $\mathrm{C}^{*}$-subalgebra. We say that $A$ admits a Lebesgue decomposition relative to $M$ if $(M_{*})|_{A}$ is norm-closed and there is a (necessarily unique) subspace $\Sigma\subset A^{*}$ such that $A^{*}=(M_{*})|_{A}\oplus_{1}\Sigma$.
It follows from [CH2025, Lemma 3.4] that $B$ itself always admits a Lebesgue decomposition relative to $M$. By contrast, there is a sense in which the existence of such a decomposition for $A$ is somewhat rare (see [CH2025] for more detail). Combining Lemma 7.3 with [CH2025, Theorem 2.4], we see that if $A$ has the F.& M. Riesz property in $B$ with respect to the convex set of restrictions to $B$ of the normal states on $M$, then $A$ admits a Lebesgue decomposition relative to $M$ that is ‘compatible’ (in a precise sense) with that of $B$. We can now give a different sufficient condition, where we don’t insist on this compatibility condition.
@@
Let $M$ be a von Neumann algebra. Let $B\subset M$ be a weak-* dense unital $\mathrm{C}^{*}$-subalgebra, and let $A\subset B$ be a unital commutative subalgebra. Let
@@
so that $\varphi\in(A^{**}(1-r_{\Gamma}))_{\perp}$. Conversely let $\varphi\in A^{*}$ such that $\varphi((1-r_{\Gamma})\cdot)=0$. By the Hahn–Banach theorem, there is $\psi\in B^{*}$ such that $\psi|_{A}=\varphi$. Since $r_{\Gamma}\in A^{\perp\perp}$, this implies that $\psi(r_{\Gamma}\cdot)|_{A}=\varphi(r_{\Gamma}\cdot)=\varphi$. Put $\psi^{\prime}=\psi(r_{\Gamma}\cdot)\in B^{*}$, which then lies in $\operatorname{AC}(\Gamma)=X$. Hence, there is $\omega\in M_{*}$ such that $\omega|_{A}=\psi^{\prime}|_{A}=\varphi$, so that $\varphi\in(M_{*})|_{A}$ as desired. ∎
@@
Let $d\geq 1$ be an integer, let ${\mathbb{B}}_{d}\subset{\mathbb{C}}_{d}$ be the open unit ball, and let ${\mathbb{S}}_{d}$ denote its topological boundary, the unit sphere. Let $\sigma$ denote surface measure on ${\mathbb{S}}_{d}$. Let $M=L^{\infty}({\mathbb{S}}_{d},\sigma),B=\mathrm{C}({\mathbb{S}}_{d})$ and let $A\subset B$ denote the ball algebra, i.e. the norm-closure of the polynomials. It is well known that the weak-* closure of $A$ in $M$ is simply $H^{\infty}({\mathbb{B}}_{d})$, by making the usual identification via radial boundary values. Further, given $f\in H^{\infty}({\mathbb{B}}_{d})$ and $0<r<1$, we define $f_{r}:\overline{{\mathbb{B}}_{d}}\to{\mathbb{C}}$ as $f_{r}(z)=f(rz)$. Then, $f_{r}\in A$, $\|f_{r}\|\leq\|f\|$ and the net $(f_{r})$ converges to $f$ in the weak-* topology of $M$. We infer that the set
@@
Next, we note that $B$ admits a Lebesgue decomposition relative to $M$ by [CH2025, Lemma 3.4]. Nevertheless, the Lebesgue decompositions of $A$ and $B$ relative to $M$ are not compatible, by [CH2025, Theorem 4.8 and Proposition 6.3]. ∎
@@
Next, recall that ${\mathcal{T}}({\mathcal{H}})$ contains the ideal of compact operators on $\mathcal{H}$}, whence it admits a Lebesgue decomposition relative to $B({\mathcal{H}})$ by [CH2025, Lemma 3.4]. Nevertheless, the Lebesgue decompositions of ${\mathcal{A}}({\mathcal{H}})$ and ${\mathcal{T}}({\mathcal{H}})$ relative to $B({\mathcal{H}})$ are not compatible, by [CH2025, Theorem 4.8 and Proposition 6.1]. ∎
We emphasize again that the incompatibility of the Lebesgue decompositions in the previous two examples shows in particular that Corollary 7.4 is applicable in some cases where [CH2025, Theorem 2.4] is not.
@@
This set will play an analogous role to the set of functionals whose restrictions to $A$ are “Henkin” (see e.g. Section 8.1 below). We now show that the F.& M. Riesz property implies a certain rigidity property for left bands containing $\Delta$. Its relation to the classical Cole-Range theorem [rudin2008, Theorem 9.6.1] will be thoroughly explained at the end of this section.
@@
    AA has the strong F.& M. Riesz property in $B$ with respect to $\Delta$.
@@
Let $\lambda$ denote arclength measure on the unit circle $\mathbb{T}$$\subset{\mathbb{C}}$. By means of radial boundary values, we may embed the algebra $H^{\infty}({\mathbb{D}})$ of bounded holomorphic functions on the open unit disc as a unital norm-closed subalgebra of $L^{\infty}({\mathbb{T}},\lambda)$. Choosing $\Delta=\{\lambda\}$, Lemma 7.2 implies that $\operatorname{AC}(\Delta)$ coincides with the weak-* continuous linear functionals on $L^{\infty}({\mathbb{T}},\lambda)$. Hence, $E(\Delta)$ coincides with the set $X$ appearing in Corollary 7.4.
Consider now the Douglas algebra $H^{\infty}({\mathbb{D}})+\mathrm{C}({\mathbb{T}})\subset L^{\infty}({\mathbb{T}},\lambda)$, which is known to be proper and norm-closed (see for instance [douglas1998, Chapter 6]). Hence, there is a non-zero bounded linear functional $\varphi$ on $L^{\infty}({\mathbb{T}},\lambda)$ that annihilates $H^{\infty}({\mathbb{D}})+\mathrm{C}({\mathbb{T}})$. Since $\mathrm{C}({\mathbb{T}})$ is weak-* dense in $L^{\infty}({\mathbb{T}},\lambda)$, we see that $\varphi$ is not weak-* continuous, so that $H^{\infty}({\mathbb{D}})$ does not have the strong F.& M. Riesz property in $L^{\infty}({\mathbb{T}},\lambda)$ with respect to $\Delta$. On the other hand, it is known that $H^{\infty}({\mathbb{D}})$ has the F.& M. Riesz property in $L^{\infty}({\mathbb{T}},\lambda)$ with respect to $\Delta$ (see [ando1978]). We may thus use Corollary 8.2 to see that the set $E(\Delta)$ is not a left band. ∎
@@ 8.1. Henkin functionals and the Cole-Range phenomenon
Let $M$ be a von Neumann algebra. Let $B\subset M$ be a unital $\mathrm{C}^{*}$-subalgebra that is weak-* dense in $M$. Let $A\subset B$ be a unital subalgebra. A bounded linear functional $\varphi:A\to{\mathbb{C}}$ is said to be Henkin relative to $M$ if it extends weak-* continuously to the weak-* closure of $A$ in $M$. This notion is in agreement with our terminology used in Section 5. Equivalently, $\varphi$ is Henkin if it extends weak-* continuously to $M$ [rudin1991FA, Theorem 3.6], so the space of Henkin functionals is simply $(M_{*})|_{A}$.
@@
Clearly, integration against $\sigma$ is in $\Omega$ by Cauchy’s formula. Furthermore, $\Omega\subset X$. Since Henkin functionals on $A$ admit a weak-* continuous extension to $M$, it follows that $X\subset A^{\perp}+\operatorname{AC}(\sigma)\subset A^{\perp}+\operatorname{AC}(\Omega)=E(\Omega)$.
@@ 9. The Amar–Lederer property
Consider $H^{\infty}({\mathbb{D}})\subset L^{\infty}({\mathbb{T}},\lambda)$ as in Example 5. Let $X$ denote the character space of $L^{\infty}({\mathbb{T}},\lambda)$. Then, via the Gelfand transform, we know that $L^{\infty}({\mathbb{T}},\lambda)\cong\mathrm{C}(X)$. Let $A\subset\mathrm{C}(X)$ denote the unital norm-closed subalgebra corresponding to $H^{\infty}({\mathbb{D}})$ under this identification. Let $\tau$ denote the state on $\mathrm{C}(X)$ corresponding to integration against $\lambda$ on $L^{\infty}({\mathbb{T}},\lambda)$. Let $\Delta=\{\tau\}$. In this case, $A$ has the Amar–Lederer property in $B$ with respect to $\Delta$ by a classical result of Amar and Lederer [AL1971]. Of course understanding this requires also the dictionary between closed sets and closed projections, and between peak sets and closed projections in $A^{\perp\perp}$, that we have discussed earlier. On the other hand, properties (iv) and (v) in Theorem 6.5 fail. To see this note that the latter properties, in the specific case of this example, imply that closed ‘Lebesgue-null’ sets are generalized peak sets. Yet [pelczynski1977, Proposition 6.3] gives a closed ‘Lebesgue-null’ $G_{\delta}$ set in this context which is not a peak set. Hence, it is not a generalized peak set, since $G_{\delta}$ generalized peak sets are peak sets [gamelin1969, Lemma II.12.1].